Which of the ordered pairs (3, 1), (0, –4), (–4, 0), (–3, –7)are solutions for the equation x – y = 4?
Answer by bucky(2189)

All points are in the standard form (x, y). So all you need to do is for each point
is to substitute the first value (which is x) for x in the equation, substitute the second
value for y in the equation, and then determine if the left side of the equation still
equals the right side of the equation.
.
Let's work them one at a time:
.
Given the point (3, 1). This means x = 3 and y = 1. Substitute these values into the equation
.
x – y = 4
.
When you substitute 3 for x and 1 for y the equation becomes:
.
3 - (1) = 4
.
This becomes:
.
3 - 1 = 4
.
which further simplifies to:
.
2 = 4
.
This is not true which means that (3, 1) is not a solution to the equation and therefore
this point is not on the graph of the given equation.
.
Next let's try the point (0, -4). This tells you to let x = 0 and y = -4 in the equation.
.
x – y = 4
.
and substituting x = 0 and y = -4 makes the equation become:
.
0 - (-4) = 4
.
This simplifies to:
.
+ 4 = 4
.
This works. Therefore (0, -4) is a solution to the equation and therefore is on the line
that is the graph for this equation.
.
Now let' try (-4, 0). This means x = -4 and y = 0. Substitute into the equation:
.
x – y = 4
.
-4 - 0 = 4
.
This does not work and so (-4, 0) is not a solution to the equation.
.
Finally, let's look at (-3, -7). Let x = -3 and y = -7:
.
x – y = 4
.
-3 - (-7) = 4
.
-3 + 7 = 4
.
4 = 4
.
This point also is a solution to the equation because both sides of the equation are equal
when the x and y values of the point are substituted into the equation.
.
In summary, there are two points in the list that are solutions to the problem.
.
Hope this helps you to understand the problem and how to work it to find the answer.